ILoveXXXTENTACION:

The scale on a model railroad train set is 3.5 millimeters to 1 foot if the length of a model car in the set is 150 millimeters whats is the approximate length of the actual car

jhonyy9:

any idea ?

ILoveXXXTENTACION:

nope.

jhonyy9:

given the scale 3,5 mm to 1 foot so you know this and you need convert this 150 mm to get the right answer do you can make it ?

jhonyy9:

ok use the rule of 3 simple 3,5 mm mean 1 foot 151 mm mean x cross multiplie and get x = (151*1)/3,5 = 151/3,5 = 43,14 ft so the aprox. length of this actuale car is 43,14 ft do you understand it ?

ILoveXXXTENTACION:

yeah, ig

jhonyy9:

ok. was my pleasure good luck bye